OpenStudy (anonymous):

-15x/x^2+8x+16 + 12/x-4 adding expressions :)

OpenStudy (anonymous):

\[\frac{-15x}{x^2}+8x+16+ \frac{12}{x-4}=\frac{(-15x)(x-4)+8x(x^2)(x-4)+16(x^2)(x-4)+12x^{2}}{(x-4)x^{2}}\] \[\frac{-15x^{2}+60x+8x^{4}-32x^{3}+16x^{3}-64x^{2}+12x{2} }{{(x-4)x^{2}}}\]\[\frac{8x^{4}-16x^{3}-67x^{2}+60x}{(x-4)x^{2}}\] \[\frac{\cancel{x}(8x^{3}-16x^{2}-67x^{1}+60)}{(x-4)x^{\cancel{2}}}\] \[\frac{(8x^{3}-16x^{2}-67x^{1}+60)}{(x-4)x}\] I can' t go more further .. regards

OpenStudy (anonymous):

the second line should be : \[\frac{-15x^{2}+60x+8x^{4}-32x^{3}+16x^{3}-64x^{2}+12x^{2} }{{(x-4)x^{2}}}\]

jhonyy9 (jhonyy9):

-15x/(x+4)^2 +12/(x-4) -the denominator common is (x+4)^2 *(x-4) so than the first term need multiplie by (x-4) and the second term by (x+4)^2 -15x(x-4) +12(x+4)^2 --------------------- = (x+4)^2 *(x-4)